(1) Consumption function: C=100+.8Y (2) Planned investment: I = 38 (3) Government spending: G=75 (4) Exports: EX=25 (5) Imports:IM= .05 Y (6) Disposable income: Y Y-T (7) Taxes: T=40 (8) Planned aggregate expenditure: AE=C+I+G+ EX-IM (9) Definition of equilibrium income: Y = AE a. What is equilibrium income in Hypothetica? What is the government deficit? What is the current account balance? b. If government spending is increased to G=80, what hap pens to equilibrium income? Explain using the government spending multiplier. What happens to imports? c. Now suppose the amount of imports is limited to IM-40 by a quota on imports. If government spending is again increased from 75 to 80, what happens to equilibrium income? Explain why the same increase in G has a bigger effect on income in the second case. What is it about the presence of imports that changes the value of the multiplier? d. If exports are fixed at EX=25, what must income be to ensure a current account balance of zero?
1
Expert's answer
2021-08-12T17:38:49-0400

(a)

"Y=EA"

"Y=C+I+G+EX-IM"

"Y=100+0.8Y+38+75+25-0.05Y"

"Y=952"


Government deficit = Government expenditure- Government income

Government expenditure ="C+I+G+EX-IM"

"=100+0.8(952)+38+75+25-0.05(952)=952"

"\therefore" government deficit= "952-40 = 912"

Current account balance = "(X-M)+NI+NT"

"=(25-47.6)+(952-40)"

"=889.4"

(b)

If government spending is increased to 80, equilibrium income is increased to "952+ 5= 957" . Equilibrium income increases in the same proportion as government spendin g.

Level of imports will be cut down by same proportion as increase in government spending.

(c)

If "IM =40"

Government spending "=80"

Equilibrium income"= 100+761.6+38+80+25-40= 965"

Equilibrium income increases from 952 to 965.


The same increase in government spending has a bigger effect on income in the second case because the level of imports has been cut down.


Value of the multiplier decreases with increase in level of imports.


(d)

Current account balance=0

"0=25-47.6+X"

"X=22.6"

Income must be 22.6 to ensure a current account balance of zero.